# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v3

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$, the largest element of which is $q$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$ and obtain generating functions for $S_{n}$ when $q=4,5,6$. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
Recently, a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences set $\mathcal{Q}$ of which the largest element is $q$ and the restricted-overlap tilings of an $(n+q)$-board (an $(n+q)\times 1$ board) with squares and $k$ combs was found. This enabled generating functions for $S^{\mathcal{Q}}_{n}$ and recursion relations for $S^{\mathcal{Q}}_{n,k}$ to be readily obtained for a variety of families of $\mathcal{Q}$ [2]. A $(w_{1},g_{1},w_{2},\ldots,g_{t-1},w_{t})$-comb is a tile composed of $t>0$ rectangular sub-tiles (called teeth) arranged in such a way that the $i$-th tooth, which has dimensions $w_{i}\times 1$, is separated from the $(i+1)$-th by a $g_{i}\times 1$ gap [2]. The length of a comb is the sum of the lengths of the teeth and gaps. Each $\mathcal{Q}$ corresponds to a length-$(q+1)$ comb $C$ (which we will refer to as a $\mathcal{Q}$-comb) defined as follows. We label the cells of $C$ from 0 to $q$. Cell 0 is always part of the first tooth. Cell $i$ (where $i=1,\ldots,q$) is part of a tooth iff $i\in\mathcal{Q}$. Note that we can always include 0 in $\mathcal{Q}$ without affecting the results (and we will do so in Theorems 16 and 17 for reasons of convenience and symmetry) since subset elements are distinct and so cannot differ by 0. This 0 element can then be regarded as corresponding to cell 0 of the comb. To ensure that there is only one way to specify the $w_{i}$ and $g_{i}$ of the $\mathcal{Q}$-comb, we insist that all teeth and gaps have positive integer length. A restricted-overlap tiling using squares and $\mathcal{Q}$-combs means that any comb can be placed so that its non-leftmost cells (i.e, cells 1 to $q$) overlap the non-leftmost cells of any other combs on the board and no other forms of overlap of tiles is allowed [2] (see Fig. 7 for examples). We let $B^{\mathcal{Q}}_{n}$ denote the number of restricted-overlap tilings of an $n$-board with $\mathcal{Q}$-combs and squares, and $B^{\mathcal{Q}}_{n,k}$ denote the number of such tilings that contain $k$ combs. The following theorem is a consequence of the bijection.
Theorem 1 (Theorem 2 in [2]).
@@
Any tiling of an $n$-board can be expressed as a tiling using metatiles which are gapless groupings of tiles that cannot be split into smaller gapless groupings by removing a gapless grouping of tiles [7]. The simplest metatiles when tiling a long enough board with squares and combs are a single square (denoted by $S$) and a $(w_{1},g_{1},\ldots,w_{t})$-comb with all the gaps filled by squares. We call the latter metatile a filled comb and it is denoted by $CS^{g}$, where $g=\sum_{i}g_{i}$. If the comb has one tooth it is just a $w_{1}$-omino (and also a metatile by itself) and $S$ and $C$ are then the only metatiles. Otherwise, as a consequence of the restricted-overlap nature of the tiling, a $C$ can always be placed so that its leftmost cell occupies the next available empty cell of the yet-to-be-completed metatile. If more empty cells within the yet-to-be-completed metatile result (which, if a $C$ is placed starting in the final cell of the final gap of the first comb on the board, will be the case iff $2w_{t}<q$ – see Lemma 2 in [2]), this can continue indefinitely and so there are an infinite number of possible metatiles.
@@
A systematic way to generate all metatiles, and thus obtain generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$, is to first construct a directed pseudograph (henceforth simply referred to as a digraph) in which each node has a label giving a bit-string representation of the remaining gaps and filled cells (a 0 representing an empty cell, a 1 a filled one) starting at the next gap in the yet-to-be-completed metatile and ending at the final filled cell [8, 2] (see Figs. 3 and 4 for examples). The 0 node represents both the empty board and the completed metatile. Each arc represents the addition of a tile and has a label showing the type of tile and a subscript giving the increase in length of the incomplete metatile resulting from placing that tile, if that increase is not 0. The length increase from adding a $C$ is given by $q+1-d$, where $d$ is the number of binary digits in the node the arc starts from but with $d=0$ for the 0 node. A walk starting and ending at the 0 node but not visiting it in between represents a metatile and is denoted by the string of $C$ and $S$ corresponding to the arcs in the order visited. For example, in Fig. 4, the loop attached to the 0 node corresponds to the $S$ metatile and the cycle starting at the 0 node and then visiting the $10^{3}$ 1 node (an abbreviation for $0001$), the $10^{2}$ 1 node, and then the 01 node before returning to the 0 node corresponds to the $CS^{2}$ (filled comb) metatile.
@@
A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 4, where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 8 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 3; Fig. 4 and Fig. 5 show digraphs that have inner cycles but no common node). For digraphs with a common node, there are simple expressions for the recursion relations for the numbers of tilings $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ in terms of lengths of various cycles in the digraph [8, 2] (we rederive these in the form of generating functions in §2). If the digraph possesses inner cycles but no common node, there is a class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [1].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for tilings with metatiles generated by a particular class of digraph (which includes the 3 inner cycle class with no common node dealt with previously [1]) without a common node. This enables us to also obtain a generating function for $S^{\{1,5\}}_{n}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
2 Generating functions for numbers of walks on digraphs with a pseudo-common node
@@
We start by reviewing further terminology concerning digraphs. For a digraph possessing a common node, a walk from the 0 node to the common node followed by a walk from the common node back to the 0 node (and so the common node is only visited once) is called a common circuit (e.g., for either choice of common node, the digraph in Fig. 3 possesses a single common circuit, namely, $C_{[5]}S^{2}$). A cycle that includes the 0 node but not the common node is called an outer cycle (e.g., in Fig. 3, the $S$ loop attached to the 0 node is the only outer cycle).
@@
If two nodes in a digraph are connected via $m-1$ intermediate nodes which are connected by walks (or sets of walks), as shown in Fig. 1b, that have associated generating functions $G_{i}(x,y)$, where $i=1,\ldots,m$, then the combined generating function for all of those walks connecting the two nodes is $\prod_{i=1}^{m}G_{i}(x,y)$.
@@
A form of the following theorem was originally stated in [8], and a more compact version re-derived in [2]. Here, using the above three lemmas, we obtain the result in terms of a bivariate generating function. We will use this theorem in §3.1 to obtain the $S^{\{1,4\}}_{n}$ generating function.
@@
The generating functions corresponding to a single traversal of the $i$-th outer cycle (represented by the loop attached to the 0 node in Fig. 2a), the $i$-th common circuit (represented by the path from the 0 node to the $\mathcal{C}$ node and back or the path $0\to\mathcal{C}\to\mathcal{R}\to 0$), and the $r$-th inner cycle ($\mathcal{C}\to\mathcal{R}\to\mathcal{C}$) are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. The $N$ inner cycles are alternative paths from the common node back to itself and so, from the parallel and loop rules, the generating function corresponding to their traversal (starting from and returning to the common node) in any order zero or more times is $1/(1-\sum_{i=1}^{N}x^{L_{r}}y^{K_{r}})$. Using the series rule, we multiply this by $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$ to obtain the generating function for the path along the $i$-th common circuit that includes an arbitrary number of trips around the inner cycles. We now apply the parallel rule to all such cases and to all outer cycles, followed by the loop rule since these can be executed any number of times. This gives
@@
The following definitions are slightly modified versions of those given in earlier work on a class of digraph with inner cycles but no common node in order to incorporate the increased complexity of the type of digraphs we are about to consider [1]. For a digraph lacking a common node, we refer to any inner cycle that can be represented as a single arc linking a node to itself as an errant loop if the digraph would only have a common node $\mathcal{P}$ if all the errant loop arcs were removed. The node $\mathcal{P}$ is then referred to as a pseudo-common node. A node with an errant loop is called an errant loop node. Evidently, a particular errant loop node $\mathcal{E}$ cannot also be a pseudo-common node; if that were the case, removing all the other errant loop arcs except the errant loop at $\mathcal{E}$ would result in a digraph with a common node implying that the loop at $\mathcal{E}$ was never an errant loop in the first place. For a digraph with at least one errant loop, a common circuit is now defined as two concatenated simple paths from the 0 node to $\mathcal{P}$ and from $\mathcal{P}$ to the 0 node. We also need to modify the definition of an outer cycle: it is now a cycle that includes the 0 node but not $\mathcal{P}$. Inner cycles (other than errant loops), outer cycles, and common circuits are said to be plain if they do not contain any errant loop nodes and non-plain otherwise.
@@
The proof is similar to that of Theorem 5; we simply have more possible types of path. The generating functions corresponding to a single traversal of the $i$-th plain outer cycle, the $i$-th plain common circuit, and the $r$-th plain inner cycle are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. On application of the series and loop rules, one sees that the generating functions corresponding to a single traversal of the $i$-th non-plain outer cycle, the $i$-th non-plain common circuit, and the $r$-th non-plain inner cycle in addition to an arbitrary number of traversals of their errant loops are, respectively, $x^{\tilde{l}_{\mathrm{o}i}}y^{\tilde{k}_{\mathrm{o}i}}/(1-x^{l_{\mathrm{oe}i}}y^{k_{\mathrm{oe}i}})$, $x^{\tilde{l}_{\mathrm{c}i}}y^{\tilde{k}_{\mathrm{c}i}}/(1-x^{l_{\mathrm{ce}i}}y^{k_{\mathrm{ce}i}})$, and $x^{L_{r}}y^{K_{r}}/(1-x^{l_{\mathrm{e}r}}y^{k_{\mathrm{e}r}})$. In Fig. 2b these are represented, respectively, by the paths $0\to\mathcal{E}_{0}\to 0$, $0\to\mathcal{P}\to\mathcal{E}_{1}\to 0$ or $0\to\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to 0$, and $\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to\mathcal{P}$, where, on reaching $\mathcal{E}_{i}$, its loop can be executed an arbitrary number of times. It should now be apparent where the terms in the denominator of (2) arise from. The first two sums are from the plain and non-plain outer cycles. The numerator of the final term is the generating function for all the common circuits (including arbitrary number of traversals of the errant loops in the case of the non-plain ones). On reaching $\mathcal{P}$ via the first part of a common circuit, the inner cycles (and their associated errant loops in the case of the non-plain ones) can be executed in any order any number of times, which is what the denominator of the final term corresponds to (by application of all three rules). ∎
@@
From Theorem 1 we have $S^{\mathcal{Q}}_{n,k}=B^{\mathcal{Q}}_{n+q,k}$, which implies that we can obtain $g^{\mathcal{Q}}(x,y)$ by removing the first $q$ terms from $G(x,y)$, namely, $1+x+x^{2}+\cdots+x^{q-1}$ (since the only way to tile a board of length less than $q$ is with no $Q$-combs), and then by dividing the result by $x^{q}$ to obtain the correct offset. ∎

### v4

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$, the largest element of which is $q$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$ and derive the generating function for $S_{n}$ when $q=4,5,6$. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
In recent work [2], which we henceforth refer to as All24, the author found a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences set $\mathcal{Q}$, of which the largest element is $q$, and the restricted-overlap tilings of an $(n+q)$-board (an $(n+q)\times 1$ board) with squares and $k$ combs. This enables generating functions for $S^{\mathcal{Q}}_{n}$ and recursion relations for $S^{\mathcal{Q}}_{n,k}$ to be readily obtained for a variety of families of $\mathcal{Q}$ [2]. A $(w_{1},g_{1},w_{2},\ldots,g_{t-1},w_{t})$-comb is a tile composed of $t>0$ rectangular sub-tiles (called teeth) arranged in such a way that the $i$-th tooth, which has dimensions $w_{i}\times 1$, is separated from the $(i+1)$-th by a $g_{i}\times 1$ gap [2]. The length of a comb is the sum of the lengths of the teeth and gaps. Each $\mathcal{Q}$ corresponds to a length-$(q+1)$ comb $C$ (which we refer to as a $\mathcal{Q}$-comb) defined as follows. We label the cells of $C$ from 0 to $q$. Cell 0 is always part of the first tooth. Cell $i$ (where $i=1,\ldots,q$) is part of a tooth iff $i\in\mathcal{Q}$. Note that we are free to include 0 in $\mathcal{Q}$ without affecting the results (and we do so in Theorems 16 and 17 for reasons of convenience and symmetry) since subset elements are distinct and so cannot differ by 0. This 0 element can then be regarded as corresponding to cell 0 of the comb. To ensure that there is only one way to specify the $w_{i}$ and $g_{i}$ of the $\mathcal{Q}$-comb, we insist that all teeth and gaps have positive integer length. A restricted-overlap tiling using squares and $\mathcal{Q}$-combs means that any comb can be placed so that its non-leftmost cells (i.e, cells 1 to $q$) overlap the non-leftmost cells of any other combs on the board and no other forms of overlap of tiles is allowed [2] (see Fig. 7 for examples). We let $B^{\mathcal{Q}}_{n}$ denote the number of restricted-overlap tilings of an $n$-board with $\mathcal{Q}$-combs and squares, and $B^{\mathcal{Q}}_{n,k}$ denote the number of such tilings that contain $k$ combs. The following theorem is a consequence of the bijection.
Theorem 1 (Theorem 2 in All24).
@@
Any tiling of an $n$-board can be expressed as a tiling using metatiles, which are gapless groupings of tiles that cannot be split into smaller gapless groupings by removing a gapless grouping of tiles [8]. The simplest metatiles when tiling a long enough board with squares and combs are a single square (denoted by $S$) and a $(w_{1},g_{1},\ldots,w_{t})$-comb with all the gaps filled by squares. We call the latter metatile a filled comb and it is denoted by $CS^{g}$, where $g=\sum_{i}g_{i}$. If the comb has one tooth it is just a $w_{1}$-omino (and also a metatile by itself), and $S$ and $C$ are then the only metatiles. Otherwise, as a consequence of the restricted-overlap nature of the tiling, a $C$ can always be placed so that its leftmost cell occupies the next available empty cell of the yet-to-be-completed metatile. If more empty cells within the yet-to-be-completed metatile result (which, if a $C$ is placed starting in the final cell of the final gap of the first comb on the board, is the case iff $2w_{t}<q$ [2, Lemma 2]), this can continue indefinitely and so there are an infinite number of possible metatiles.
@@
A systematic way to generate all metatiles, and thus obtain generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$, is to first construct a directed pseudograph (henceforth simply referred to as a digraph) in which each node has a label giving a bit-string representation of the remaining gaps and filled cells (a 0 representing an empty cell, a 1 a filled one) starting at the next gap in the yet-to-be-completed metatile and ending at the final filled cell [9, 2] (see Figs. 3, 4, and 5 for examples). The 0 node represents both the empty board and the completed metatile. Each arc represents the addition of a tile and has a label showing the type of tile and a subscript giving the increase in length of the incomplete metatile resulting from adding that tile, if that increase is not 0. The length increase from adding a $C$ is given by $q+1-d$, where $d$ is the number of binary digits in the node the arc starts from, but with $d=0$ for the 0 node. A walk starting and ending at the 0 node, but not visiting it in between, represents a metatile and is denoted by the string of $C$ and $S$ corresponding to the arcs in the order visited. For example, in Fig. 4, the loop attached to the 0 node corresponds to the $S$ metatile, and the cycle starting at the 0 node and then visiting the $10^{3}$ 1 node (an abbreviation for $0001$), the $10^{2}$ 1 node, and then the 01 node before returning to the 0 node corresponds to the $CS^{2}$ (filled comb) metatile.
@@
A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 4, where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 8 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 3; Fig. 4 and Fig. 5 show digraphs that have inner cycles but no common node). For digraphs with a common node, there are simple expressions for the recursion relations for the numbers of tilings $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ in terms of lengths of various cycles in the digraph [9, 2] (we rederive these in the form of generating functions in §2). If the digraph possesses inner cycles, but no common node, there is a class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [1].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ shows that there are various connections between some types of restricted combinations and some types of restricted permutations. One also finds relations to compositions and bit strings. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for tilings with metatiles generated by a particular class of digraph (which includes the 3 inner cycle class with no common node dealt with previously [1]) without a common node. This enables us to easily obtain a generating function for $S^{\{1,5\}}_{n}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. We point out an elementary connection between all instances of $\mathcal{Q}$ and bit strings in §5. This forms the basis for an algorithm for efficiently calculating $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$ for any given $\mathcal{Q}$, as detailed via a heavily annotated C program listed in the Appendix. In §5, we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain. In the discussion (§6), we mention transfer matrices as an alternative way to obtain the $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ generating functions from digraphs.
2 Generating functions for the numbers of walks on digraphs
@@
We start by reviewing further terminology concerning digraphs. For a digraph possessing a common node, a path from the 0 node to the common node followed by a path from the common node back to the 0 node (and so the common node is only visited once) is called a common circuit (e.g., for either choice of common node, the digraph in Fig. 3 possesses a single common circuit, namely, $C_{[5]}S^{2}$). A cycle that includes the 0 node, but not the common node, is called an outer cycle (e.g., in Fig. 3, the $S$ loop attached to the 0 node is the only outer cycle).
@@
If two nodes in a digraph are connected via $m-1$ intermediate nodes, which are connected by walks (or sets of walks), as shown in Fig. 1b, that have associated generating functions $G_{i}(x,y)$, where $i=1,\ldots,m$, then the combined generating function for all of those walks connecting the two nodes is $\prod_{i=1}^{m}G_{i}(x,y)$.
@@
A form of the following theorem was originally stated by Edwards and Allen [9], and a more compact version re-derived later [2]. Here, using the above three lemmas, we obtain the result in terms of a bivariate generating function. We use this theorem in §3.1 to obtain the $S^{\{1,4\}}_{n}$ generating function.
@@
The generating functions corresponding to a single traversal of the $i$-th outer cycle (represented by the loop attached to the 0 node in Fig. 2a), the $i$-th common circuit (represented by the path from the 0 node to the $\mathcal{C}$ node and back or the path $0\to\mathcal{C}\to\mathcal{R}\to 0$), and the $r$-th inner cycle ($\mathcal{C}\to\mathcal{R}\to\mathcal{C}$) are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. The $N$ inner cycles are alternative paths from the common node back to itself and so, from the parallel and loop rules, the generating function corresponding to their traversal (starting from and returning to the common node) in any order zero or more times is $1/(1-\sum_{i=1}^{N}x^{L_{r}}y^{K_{r}})$. Using the series rule, we multiply this by $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$ to obtain the generating function for the walk that consists of the paths to and from $\mathcal{C}$ along the $i$-th common circuit and an arbitrary number of trips around the inner cycles. We now apply the parallel rule to all such cases and to all outer cycles, followed by the loop rule since these can be executed any number of times. This gives
@@
The following definitions are slightly modified versions of those given in earlier work on a class of digraph with inner cycles but no common node in order to incorporate the increased complexity of the type of digraphs we are about to consider [1]. For a digraph lacking a common node, we refer to any inner cycle that can be represented as a single arc linking a node to itself as an errant loop if the digraph would only have a common node $\mathcal{P}$ if all the errant loop arcs were removed. The node $\mathcal{P}$ is then referred to as a pseudo-common node. A node with an errant loop is called an errant loop node. Evidently, a particular errant loop node $\mathcal{E}$ cannot also be a pseudo-common node; if that were the case, removing all the other errant loop arcs except the errant loop at $\mathcal{E}$ would result in a digraph with a common node, implying that the loop at $\mathcal{E}$ was never an errant loop in the first place. For a digraph with at least one errant loop, a common circuit is now defined as two concatenated simple paths from the 0 node to $\mathcal{P}$ and from $\mathcal{P}$ to the 0 node. We also need to modify the definition of an outer cycle: it is now a cycle that includes the 0 node but not $\mathcal{P}$. Inner cycles (other than errant loops), outer cycles, and common circuits are said to be plain if they do not contain any errant loop nodes and non-plain otherwise.
@@
The proof is similar to that of Theorem 5; we simply have more possible types of walk. The generating functions corresponding to a single traversal of the $i$-th plain outer cycle, the $i$-th plain common circuit, and the $r$-th plain inner cycle are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. On application of the series and loop rules, one sees that the generating functions corresponding to a single traversal of the $i$-th non-plain outer cycle, the $i$-th non-plain common circuit, and the $r$-th non-plain inner cycle in addition to an arbitrary number of traversals of their errant loops are, respectively, $x^{\tilde{l}_{\mathrm{o}i}}y^{\tilde{k}_{\mathrm{o}i}}/(1-x^{l_{\mathrm{oe}i}}y^{k_{\mathrm{oe}i}})$, $x^{\tilde{l}_{\mathrm{c}i}}y^{\tilde{k}_{\mathrm{c}i}}/(1-x^{l_{\mathrm{ce}i}}y^{k_{\mathrm{ce}i}})$, and $x^{L_{r}}y^{K_{r}}/(1-x^{l_{\mathrm{e}r}}y^{k_{\mathrm{e}r}})$. In Fig. 2b these are represented, respectively, by the walks $0\to\mathcal{E}_{0}\to 0$, $0\to\mathcal{P}\to\mathcal{E}_{1}\to 0$ or $0\to\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to 0$, and $\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to\mathcal{P}$, where, on reaching $\mathcal{E}_{i}$, its loop can be executed an arbitrary number of times. It should now be apparent where the terms in the denominator of (2) arise from. The first two sums are from the plain and non-plain outer cycles. The numerator of the final term is the generating function for all the common circuits (including arbitrary number of traversals of the errant loops in the case of the non-plain ones). On reaching $\mathcal{P}$ via the first part of a common circuit, the inner cycles (and their associated errant loops in the case of the non-plain ones) can be executed in any order any number of times, which is what the denominator of the final term corresponds to (by application of all three rules). ∎
@@
From Theorem 1 we have $S^{\mathcal{Q}}_{n,k}=B^{\mathcal{Q}}_{n+q,k}$, which implies that we can obtain $g^{\mathcal{Q}}(x,y)$ by removing the first $q$ terms from $G(x,y)$, namely, $1+x+x^{2}+\cdots+x^{q-1}$ (since the only way to tile a board of length less than $q$ is with no $\mathcal{Q}$-combs), and then by dividing the result by $x^{q}$ to obtain the correct offset. ∎